OpenStudy (anonymous):

The probability of at least three successes in six trials of a binomial experiment in which p = 1/2

OpenStudy (ybarrap):

http://openstudy.com/study#/updates/53127432e4b0966c639ba915 $$ \large{ P(X \ge3)=1-P(X<3)\\ =1-P(X\le2)\\ =1-(P(X=0)+P(X=1)+P(X=2))\\ =1-\sum_{i=0}^2 \binom{6}{i}p^iq^{6-i}\\ =1-\sum_{i=0}^2 \binom{6}{i}0.5^6\\ =1-0.5^6\sum_{i=0}^2 \binom{6}{i}\\ =1-0.5^6\left(\binom{6}{0}+\binom{6}{1}+\binom{6}{2}\right )\\ =1-0.5^6\left(1+6+15\right )\\ =1-0.5^6\left(32\right )\\ =1-0.5^6\left(2^5\right )\\ =1-\cfrac{2^5}{2^6}\\ =1-\cfrac{1}{2}\\ =\cfrac{1}{2} } $$